The length of a rectangle is the width minus 2 units. The area of the rectangle is 8 units. What is the length, in units, of the rectangle?
Answer by Cromlix(4381)

Hi there,
Make width = x
Length = x - 2
Area of rectangle
= length x width
8 = (x - 2)x
x^2 - 2x - 8 = 0
(x - 4)(x + 2) = 0
x + 2 = 0
x = -2 (Disregard as -ve value)
x - 4 = 0
x = 4
Width = 4 units
Length = 2 units
Hope this helps :-)